How can I find correlation coefficient decision point on Excel?

I'm looking for when n=230
Answer #1

here you first find critical value t at level of significance ; suppose here level of sigifcance is 0.05 and degree of freedom

=n-2=230-2=228

hence critical value t =tinv(0.05,228) =1.970   (excel formula is in bold)

therfore crtiical correlation coefficient =sqrt(t2/(t2+n-2)) =sqrt(1.9702/(1.970+230-2))=0.129

( please revert for any clarification required)